# M3D: Manifold-based Domain Adaptation with Dynamic Distribution for Non-Deep Transfer Learning in Cross-subject and Cross-session EEG-based Emotion Recognition

### v2

@@ V-A Ablation Study
F1-Score 69.68 72.56 71.28 76.11
V-B Investigating factors affecting model performance in different modules
To evaluate the efficacy of TCA in the manifold feature transformation, we assess the model performance by comparing the use of TCA against substituting TCA with Principal Component Analysis (PCA). As shown in Table X, the model exhibits significantly better performance in terms of accuracy and F1-Score when comparing TCA with PCA. This improvement can be attributed to TCA’s ability to enhance model performance by simultaneously mapping and reducing dimensions, and by maximizing the similarity between datasets from two domains. It would facilitate more effective dynamic distribution alignment and classifier categorization.
TABLE XII: Performance comparison with different ensemble learning.
Methods Last Averaging Voting LinkCluE
@@ V-B Investigating factors affecting model performance in different modules
F1-Score 69.94 66.97 71.59 76.11
TABLE XIII: The emotion recognition accuracy (%) without and with ensemble learning in iterative learning process.
Iteration 2 5 10 20 30 50
w/o ensemble learning 72.85 75.09 74.61 74.09 74.37 73.45
with ensemble leaning 80.39 83.81 84.57 83.84 83.34 82.55
@@
To evaluate the effectiveness of dynamic distribution alignment, we opted for a fixed $\mu$ value approach, rather than using an adaptive $\mu$. The $\mu$ values are set at 00, $0.5$, and $1$. At $\mu=0$, the model exclusively focuses on the marginal distribution. With $\mu=0.5$, both marginal and conditional distributions are considered equally. At $\mu=1$, the model solely emphasizes the conditional distribution. As reported in Table XI, it shows that employing an adaptive $\mu$ can significantly enhance model performance, resulting in a 4.26% improvement in accuracy. The results demonstrate that dynamic distribution alignment can effectively quantify the disparity between source and target domains and accurately evaluate the significance of the differences between marginal and conditional distributions. This capability is particularly advantageous for the classification performance, which is tailored for addressing variations between distributions.
@@
To evaluate the impact of ensemble learning on model performance, we adjust our approach by incorporating various ensemble learning techniques, namely the averaging method, the voting method, and the LinkCLuE method. Note here that LinkCLuE is the method implemented in the proposed M3D. Based on the obtained classification results in the classifier learning ($\hat{y}_{t}^{(\iota)}$ ($\iota=1,\cdots,10$)), the effectiveness of ensemble learning is analyzed. Furthermore, we drew comparisons to scenarios devoid of ensemble learning strategies by directly employing the final classification outcome from the last iteration loop in the classifier learning, denoted as $\hat{y}_{t}^{(\iota)}$ (with $\iota=10$), which serves as our baseline. The experimental comparison results are reported in Table XII. It shows that ensemble learning with voting and LinkCluE could enhance the reliability of classification results and improve the overall performance. This confirms the effectiveness of ensemble algorithms. Furthermore, the LinkCLuE method outperforms the voting method, demonstrating a superior ability to synthesize classification outcomes across iterations and to discern the underlying patterns of the classification results.
V-C Parameter effect
To examine the classifier learning process, we also investigate how sensitive the model is to changes in the number of iterations ($\iota$) in the learning process. Here, we vary $\iota$ value in a range of $[2,5,10,20,30,50]$, and present the corresponding learning results in Table XIII. It shows the model’s classification performance experiences fluctuations during the initial few iterations but achieves stability beyond $\iota=10$ iterations. This demonstrates the training advantage of M3D in cross-domain tasks.
VI Conclusion
In this paper, we propose a novel non-deep transfer learning framework, termed M3D, for cross-subject and cross-session EEG-based emotion recognition, as well as clinical mental disorder recognition. The framework employs a manifold-based domain adaptation with dynamic distribution alignment, adaptively balancing marginal and conditional distributions in the Grassmann manifold space. This enhances intrinsic data representation and reduces domain disparity. Additionally, optimized classifier learning with ensemble techniques improves robustness and reliability in handling EEG data variations. Extensive experiments on two benchmark databases validate the framework under two evaluation protocols. Furthermore, we demonstrate the model’s effectiveness on a real clinical MDD dataset. Compared to existing methods, M3D achieves promising results against non-deep approaches and performance comparable to deep learning models, with an average improvement of 4.47% in classification accuracy. These findings highlight the potential of non-deep transfer learning in addressing individual and session variations in affective brain-computer interfaces. On the other hand, in the dynamic distribution alignment module, we did not impose constraints on the adaptive factor $\mu$ during the iteration process, which may lead to suboptimal feature alignment in certain cases. Addressing this limitation in future work could further enhance the stability and generalizability of the framework.

### v3

@@ 9 The effect of the parameter
Sensitivity 70.89 75.42 71.69 76.04
Specificity 82.80 84.60 84.96 88.78
Precision 69.38 70.94 72.03 76.94
F1-Score 70.13 73.11 71.86 76.49
AUROC 84.12 86.16 85.31 88.36
NPV 83.80 87.35 84.74 88.27
Ensemble Methods Last Averaging Voting LinkCluE
@@
Sensitivity 74.37 79.58 73.38 76.04
Specificity 82.21 78.23 84.38 88.78
Precision 66.86 57.86 70.94 76.94
F1-Score 70.41 67.00 72.14 76.49
AUROC 84.68 84.17 85.50 88.36
NPV 86.92 91.07 85.92 88.27
Figure 15: The mean accuracies (%) of the proposed M3D under different sample sizes.
10 Theoretical analysis of the computational complexity
We conduct a comprehensive theoretical analysis of the computational complexity of the proposed M3D framework, taking into account its main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning.
Let $n$ and $m$ be the sample sizes of the source and target domains respectively, and $d$ be the dimensionality of the reduced features. In the dynamic distribution alignment module, the computational cost is mainly consumed in the dimensionality reduction process (such as Transfer Component Analysis, TCA) and the computation of the geodesic kernel $G$ on the Grassmann manifold. Specifically, the time complexity of the dimensionality reduction is $O(d(n+m)^{2})$. Calculating the geodesic kernel $G$ involves matrix operations such as singular value decomposition (SVD), leading to a time complexity of $O(d^{3})$. Key computations in the dynamic distribution alignment module include calculating the adaptive factor $\mu$ and constructing the marginal and conditional distribution matrices. The calculation of $\mu$ incurs a complexity of $O(n+m)$, while computing the distribution matrices requires $O(C(n+m)^{2})$, where $C$ is the total number of classes. In the classifier learning module, it involves solving an optimization problem through matrix inversion, which leads to a time complexity of $O((n+m)^{3})$. Since this process is iteratively performed $l$ times, the overall complexity of this module is $O(l(n+m+C(n+m)^{2}+(n+m)^{3}))$. In the ensemble learning module, the computational cost is dominated by calculating the similarity matrix with $O(m^{2})$ complexity, followed by hierarchical clustering with $O(m^{2}\log m)$ complexity.
By combining the complexities of all modules, the overall time complexity of M3D can be expressed as:
$O\bigl{(}d(n+m)^{2}+d^{3}+l(n+m+C(n+m)^{2}\\ +(n+m)^{3}+m^{2}+m^{2}\log m\bigr{)}.$ (31)
In our experimental setting, considering that $l$, $d$, $C$ and $m$ are relatively small compared to $(n+m)$, the dominant term is $O(l(n+m)^{3})$. This shows that the proposed M3D maintains a relatively low computational complexity compared to deep learning models, which typically involve repeated backpropagation and large-scale parameter optimization.
11 Investigating factors affecting model performance
We also conduct a series of experiments to study the factors that affect the performance of the model. To evaluate the efficacy of the Grassmann manifold, we conduct a comparative experiment by replacing it with the Stiefel manifold in the manifold feature transformation module. The Grassmann manifold excels at depicting the geometric structure of subspaces in high-dimensional data, precisely capturing the internal connection between EEG signals and emotions, while the Stiefel manifold mainly focuses on the manifold composed of orthogonal matrices and lacks the specific ability to extract EEG emotion features. As shown in Table 13, after the replacement, the accuracy drops to 67.94%, and the average performance of 7 indicators decreases by 16.14%. This further demonstrates that the Grassmann manifold has significant advantages in extracting EEG features related to emotions, which is more conducive to improving the model’s emotion recognition performance.
To evaluate the efficacy of TCA in the manifold feature transformation, we assess the model performance by comparing the use of TCA against substituting TCA with Principal Component Analysis (PCA). As shown in Table 13, the model exhibits significantly better performance in terms of accuracy and F1-Score when comparing TCA with PCA. This improvement can be attributed to TCA’s ability to enhance model performance by simultaneously mapping and reducing dimensions, and by maximizing the similarity between datasets from two domains. It would facilitate more effective dynamic distribution alignment and classifier categorization.
@@ 11 Investigating factors affecting model performance
To evaluate the effectiveness of dynamic distribution alignment, we opt for a fixed $\mu$ value approach, rather than using an adaptive $\mu$. The $\mu$ values are set at 00, $0.5$, and $1$. At $\mu=0$, the model exclusively focuses on the marginal distribution. With $\mu=0.5$, both marginal and conditional distributions are considered equally. At $\mu=1$, the model solely emphasizes the conditional distribution. As reported in Table 13, it shows that employing an adaptive $\mu$ can significantly enhance model performance, resulting in a 4.26% improvement in accuracy. The results demonstrate that dynamic distribution alignment can effectively quantify the disparity between source and target domains and accurately evaluate the significance of the differences between marginal and conditional distributions. This capability is particularly advantageous for the classification performance, which is tailored for addressing variations between distributions.
@@
To evaluate the impact of ensemble learning on model performance, we adjust our approach by incorporating various ensemble learning techniques, namely the averaging method, the voting method, and the LinkCluE method. Note here that LinkCluE is the method implemented in the proposed M3D. Based on the obtained classification results in the classifier learning ($\hat{y}_{t}^{(\iota)}$ ($\iota=1,\cdots,10$)), the effectiveness of ensemble learning is analyzed. Furthermore, we draw comparisons to scenarios devoid of ensemble learning strategies by directly employing the final classification outcome from the last iteration loop in the classifier learning, denoted as $\hat{y}_{t}^{(\iota)}$ (with $\iota=10$), which serves as our baseline. The experimental comparison results are reported in Table 13. It shows that ensemble learning with voting and LinkCluE could enhance the reliability of classification results and improve the overall performance. This confirms the effectiveness of ensemble algorithms. Furthermore, the LinkCluE method outperforms the voting method, demonstrating a superior ability to synthesize classification outcomes across iterations and to discern the underlying patterns of the classification results.
12 Visualization of Feature Alignment
We employ t-SNE visualizations and EEG topographic maps to intuitively demonstrate the alignment of source and target features before and after adaptation. Specifically, we utilize the t-SNE algorithm [63] to visualize the source and target features from the SEED database both before and after alignment. The results are illustrated in Fig. 16. To enhance clarity, we assign different colors to the samples with different emotions and assign different shapes to distinguish samples from different domains. The results demonstrate that M3D reduces the distribution discrepancy among the two domains throughout the training process. And it progressively separates the samples with different emotions to minimize the emotion classification error. In summary, the results provide supplementary evidence to support the efficacy of the proposed M3D in EEG-based emotion.
In the topographic analysis, we identify the important brain patterns for emotion recognition by computing the mutual information between the brain patterns and prediction labels. The EEG topographic maps can intuitively display the spatial distribution characteristics of neural activities related to different emotions. Specifically, we identify the important brain patterns for emotion recognition by computing the mutual information between the brain patterns and prediction labels. We have the data input of the target domain, given as $X_{t}$, which is a $M\times 310$ matrix. Here, $M$ is the sample size in the target domain, and 310 refers to the extracted DE features from 62 channels at 5 frequency bands. The corresponding model prediction result is $\hat{Y_{t}}$, with a size of $M\times 3$. The columns in $\hat{Y_{t}}$ indicate the prediction probabilities of different emotions (negative, neutral, and positive). Then, we estimate the mutual information between the input features $X_{t}$ and and the prediction result $\hat{Y_{t}}$ using the non-parametric method as stated in [64, 65, 66]. The obtained mutual information matrix is termed as $I(X_{t},\hat{Y_{t}})\in{R^{3\times 310}}$, indicating a quantification of the inherent dependence between the EEG patterns and the model prediction results. $I(X_{t},\hat{Y_{t}})$ is further normalized to $[0,1]$, and a larger value refers to a greater informativeness of the EEG patterns to the model prediction. For the $\hat{Y_{t}}$ before alignment, it is the prediction result obtained by directly inputting the DE features into the DT classifier. While for the $\hat{Y_{t}}$ after alignment, it is the final result obtained by inputting the DE features into the model M3D.
Figure 16: A visualization of the learned feature representations (a) before alignment, (b) after alignment. Here, the circle and triangle represent the source domain ($S$) and the target domain ($T$). The red, blue and green colors indicate happy, sad and neural emotions.